OpenStudy (anonymous):

For two random variables X and Y, if E(X)=5, E(X^2)=28, E(Y)=6, E(Y^2)=40, E(XY)=24. What is the value of Var(2X-Y)?

OpenStudy (kirbykirby):

\(Var(2X-Y)=2^2Var(X)+Var(Y)+2(-2)Cov(X,Y)\) \(Var(X)=E(X^2)=[E(X)]^2\) \(Cov(X,Y)=E(XY)-E(X)E(Y)\)

OpenStudy (kirbykirby):

\(Var(X)=E(X^2)-[E(X)]^2\) ** sorry

OpenStudy (anonymous):

Thank you!

OpenStudy (kirbykirby):

=] yw